# Can a Quantum Computer Simulate Nuclear Magnetic Resonance Spectra Better than a Classical One?

### v1

@@ IV Exactly Solvable Examples
To lend credence to the utility of this metric as a means of comparing spectral functions, and to give an intuitive sense for what different values of this metric mean, we display in Figure 3 the spectral function of the molecule Limonene at low broadening and low field, computed for several different maximum cluster sizes, in each case comparing against the exact result (which in this case corresponds to a molecule with sixteen nuclear spins). In these plots, rather than plotting against the angular frequency directly, we plot against the quantity
$\Delta=\frac{\omega-\omega_{\text{ref}}}{\omega_{\text{ref}}},$ (12)
where $\omega_{\text{ref}}$ is the Larmor frequency of protons. Note, in experiments the reference frequency is typically set by the resonance of the protons in tetramethylsilane. This shifted and scaled variable has peaks located at the values of the chemical shifts $\delta$, rather than the Larmor frequencies themselves, which makes the identification of chemical groups simpler, and is the convention in the NMR literature (this is sometimes referred to as “plotting against the chemical shift”).
@@
With this understanding of the error metric in mind, we present results for the convergence of three selected molecules as a function of maximum cluster size, all of which contain sixteen nuclear spins. We display the high broadening case in Figure 1 and the low broadening case in Figure 2. As we would expect, the convergence generally improves for higher field and higher broadening, with the broadening seemingly having the more pronounced effect on accuracy. It is true that the rate of convergence of the three molecules is not necessarily identical. However, we highlight the fact that there are no sudden jumps in accuracy, either up or down, as a function of maximum cluster size, and that the error of the approximation is (almost always) decreasing. This fact will be important for attempting to benchmark the performance of the solver when applied to molecules for which we do not have an exact solution.
In Appendix D we provide convergence data for additional selected molecules, which all display the same general behaviour.
@@
Here we discuss the performance of the solver when applied to larger molecules for which we do not have an exact solution. Instead of comparing against an exact case, the solution for each cluster size is compared against the largest size we can simulate (referred to as the “reference” size). As we examine the spectral function corresponding to smaller cluster sizes, if we eventually reach a cluster size beyond which we no longer see any significant differences between the smaller size and the reference size, then we will consider the solutions to have converged. While this does not represent a rigorous proof of convergence, it is nevertheless plausible, given that we have not found any examples in the exactly solvable case in which the accuracy of the solution suddenly jumped, especially not after reaching a plateau.
@@
In Figures 6 and 7 we display the results of our analysis for a selection of ten molecules of various size, all of them with a reference cluster size of seventeen. Once again, we see that the convergence is better for higher field and higher broadening, and the agreement between a cluster size of sixteen and a cluster size of seventeen is exceptionally good in all cases. The worst convergence is for the case of Friedelin at low broadening, with error values of -2.77, -2.70, and -2.75 for high, low, and very low field, respectively. We also notice that the error values for the molecule tert-Butyl(tert-butylphosphanyl)phosphane, also named 1,2-di-tert-butyl-diphosphane, (referred to here as “Diphosphane”) seem to progress in a somewhat erratic fashion, for several parameter regimes. We briefly investigate the spectral functions of both of these molecules.
IV.1 Friedelin
@@
In Figure 9 we display the spectral function of Friedelin, for the case of low field and low broadening, which represents the worst convergence of any spectral function. While there are some small details which are not perfectly captured, the majority of the larger features appear to be resolved, to the extent that it is not clear whether a more precise simulation would add to the practical interpretation of the NMR spectrum. For the case of Friedelin, this implies that our solver is capable of reaching the level of accuracy necessary for industrial applications, especially since this level of broadening is already somewhat experimentally unrealistic. However, the lack of perfect agreement suggest that there may still exist some room for possible improvement from the use of a quantum device, if not for Friedelin, then for another, similarly large molecule.
IV.2 Diphosphane
@@
Figure 11: The spectral function for Diphosphane, for the case of high field and low broadening.
@@
To understand the origin of this behaviour, in Figure 10 we show a schematic representation of the structure of the Diphosphane molecule. This molecule consists of two central Phosphorous atoms each coupled to three Methyl groups, which together contain a total of nine Hydrogen atoms. Additionally, two Hydrogens are strongly coupled to the two 13P nuclei. As these couplings are direct and not mediated via the carbon atomes, the molecule features unusually high spin-spin interaction. Furthermore, the gyromagnetic ratio of the Phosphorous nucleus is roughly half that of the Hydrogen nucleus, implying that there are two very different energy scales at play in this system. Given these unusual features of the Diphosphane molecule, it is not surprising that a clustering approach may produce inaccurate results.
@@
Figure 12: The spectral function for Diphosphane, for the case of high field and low broadening, at larger cluster sizes, which we can obtain due to the symmetry of the molecule.
@@
To underline this conclusion, we can in fact compare to the exact solution for Diphospane, due to the high level of symmetry exhibited by this molecule. The two phosphorous atoms have the same chemical shift as each other, which is also true for the two central Hydrogen atoms. These two central Hydrogen atoms couple with the same strength to their respective phosphorous atoms. However, much more significant is the fact that all of the eighteen Hydrogen atoms contained in the Methyl groups have identical chemical shifts, and also couple identically to their respective phosphorous atoms. Thus, all of the “magnetically equivalent” spins can be combined into a single (higher representation) spin object, by exploiting the local SU(2) symmetry. Specifically, this reduction leads to an effective system of two spin-9/2 particles coupled to four spin-1/2 particles, which can be solved in approximately 10 minutes of computational time on a laptop with 32 Gigabytes of memory and an Intel(R) Core(TM) Ultra 7 165U processor.
@@
In Figure 12 we again display the spectral function of Diphosphane at the two relevant portions of the spectrum, this time for larger cluster sizes. We continue to see the oscillating behaviour at higher chemical shifts, right up to the exact solution. In addition, we see a dramatic qualitative change in the spectrum at lower chemical shift as we increase the cluster size from 21 to 22 spins (the full size of the system). It is therefore indeed true that a clustering approach fails for such a molecule. However, it should be stressed that our method of solving for the exact spectrum of Diphosphane (which involves combining magnetically equivalent spins) only works well due to the high level of symmetry in this molecule, but for similar molecules with some level of disorder in their shifts and couplings, this specific approach would not be successful.
V Conclusion
@@
In this work we have presented results for the performance of our NMR solver when applied to a collection of selected molecules. In almost every case, we have found good agreement for every parameter regime of practical relevance, with the performance increasing with higher magnetic fields and higher broadening. This demonstrates the utility of our clustering-based solver as a powerful tool for understanding molecular NMR spectra, even for molecules that we can not solve exactly.
The only case in which we see poor performance is the case of a phosphorous-containing molecule exhibiting unusually strong spin-spin interactions, due to the direct coupling of certain Hydrogen atoms to the 13P nuclei. While for this specific molecule we were able to calculate the exact spectrum due to its high level of symmetry, there may in fact exist molecules with similar (but not identical) structure, where this would no longer be possible. Such considerations point towards a possible first line of investigation in the direction of quantum advantage.
@@
In general, the convergence of our solver becomes slower for decreasing broadening or lower magnetic fields. Is it then experimentally realistic to consider reaching even lower broadening values and magnetic fields? The answer to this question can be found in the field of zero field NMR spectroscopy [58], in which one makes use of polarized samples to measure the NMR spectrum at magnetic field strengths of only a few nanotesla, such that the Zeemann term can then essentially be neglected. Interestingly, the line width in these experiments can be on the order of 0.01 Hz. We therefore expect classical solvers (including our own) to struggle in this regime, leading to a second path towards potential quantum advantage [35].
We are actively investigating both of these approaches, and hope to report on them in more depth in the near future.
@@ Appendix A Derivation of the Spectral Function
$C\left(\omega\right)\propto\eta\sum_{n,m}\frac{\langle E_{n}|M^{-}|E_{m}\rangle\langle E_{m}|M^{+}|E_{n}\rangle}{\eta^{2}+[\omega-(E_{n}-E_{m})]^{2}}$ (13)
is connected to measurements made in an NMR experiment, specifically the so-called pulse-acquire-experiment. To this end, we will first discuss what properties are measured in such an experiment, and then derive from them the spectral function.
@@ A.1 The pulse acquire experiment
In an NMR experiment, the magnetic field resulting from the magnetization of the sample is measured. The magnetic moment of a single molecule in the sample along a specified direction is given by the expectation value of the appropriate spin operator,
$m^{\alpha}=\hbar\sum_{i}\gamma_{i}\langle\hat{I}^{\alpha}\rangle=\hbar\langle M^{\alpha}\rangle$ (14)
@@
$M^{\alpha}\equiv\sum_{i}\gamma_{i}\hat{I}^{\alpha},$ (15)
with $\alpha\in\{x,y,z\}$. As the induced magnetic field is proportional to the magnetic moment, it is sufficient to add up the contributions from the magnetic moments of all of the molecules in the sample, and multiply by the appropriate permeability to find the total magnetic field. If we assume that the sample is reasonably uniform, then the magnetic field will also be uniform, and will simply be some constant multiple of the magnetic moment of one molecule. Since we are not interested in the overall normalization of the spectral function, we can simply study this individual magnetic moment. Our task then is to understand how the quantity above evolves in time in an NMR experiment.
@@
After a molecule is placed in an NMR spectrometer, it is allowed to come to thermal equilibrium, a state described by the density matrix
$\rho=\frac{1}{Z}e^{-\beta\hbar H}.$ (16)
@@
with $Z$ the partition sum. For a spectrometer with a natural frequency $\nu$ of 500 MHz, the characteristic energy scale of the molecule is
$h\nu\approx 3.3\times 10^{-25}~\text{J},$ (17)
while the thermal energy scale is
$k_{B}T\approx 4.1\times 10^{-21}~\text{J}.$ (18)
@@
Since the thermal energy scale is several orders of magnitude larger than the characteristic energy scale of the molecule, it is an incredibly good approximation to take
@@
$\begin{split}Z\rho&=I-\beta\hbar H\\ &=I-\beta\hbar\left[-\sum_{l}\gamma_{l}B^{z}\hat{I}^{z}_{l}+2\pi\sum_{k<l}J_{kl}\mathbf{\hat{I}}_{k}\cdot\mathbf{\hat{I}}_{l}\right],\end{split}$ (19)
@@
where we have again ignored the correction due to the chemical shifts. Since the coupling terms are several orders of magnitude smaller than the Zeeman terms, the expectation values of the two-site operators are negligible, and we can simply take
$Z\rho=I+\beta\hbar B^{z}\sum_{l}\gamma_{l}\hat{I}^{z}_{l}=I+\beta\hbar B^{z}M^{z}$ (20)
The sample is then subjected to a sequence of magnetic pulses, which are engineered to induce a $\pi/2$ rotation along the X axis for each active nuclei present in the sample, the dynamics of which we take to be effectively instantaneous, so that the density matrix evolves according to
$\rho\to\exp\left(-i\frac{\pi}{2}\hat{I}^{x}\right)\rho\exp\left(+i\frac{\pi}{2}\hat{I}^{x}\right)$ (21)
where we have defined
$\hat{I}^{x}=\sum_{i}\hat{I}^{x}_{i}\,.$ (22)
and restricted the result to spin-1/2 systems. The extension to general spin-S systems is straightforward. Applying this transformation to the thermal density matrix, we have
$Z\rho_{0}\to I-\beta\hbar B^{z}M^{y}$ (23)
as the initial density matrix after the pulse.
After this pulse, the density matrix of the molecule will evolve in time according to
$\rho\left(t\right)=e^{-iHt}\rho_{0}e^{+iHt}\propto I-\beta\hbar B^{z}e^{-iHt}M^{y}e^{+iHt}.$ (24)
@@
$m^{\alpha}=\hbar\langle M^{\alpha}\rangle=\hbar\text{Tr}\left[M^{\alpha}\rho\left(t\right)\right].$ (25)
Since $M^{\alpha}$ is traceless, only the second term in the density matrix contributes to the trace, and we have
$m^{x}=\hbar\langle M^{x}\rangle\propto\text{Tr}\left[M^{x}e^{-iHt}M^{y}e^{+iHt}\right],$ (26)
@@
$m^{y}=\hbar\langle M^{x}\rangle\propto\text{Tr}\left[M^{y}e^{-iHt}M^{y}e^{+iHt}\right],$ (27)
@@
$m^{\alpha}\to e^{-\eta t}m^{\alpha}\,.$ (28)
Note that in the derivation we have made use of a high temperature and high magnetic field expansion. However, due to the small energy scale of the NMR system, it is valid for temperatures above 1K and magnetic fields larger than the Earth’s magnetic field.

### v2

@@ IV Exactly Solvable Examples
To lend credence to the utility of this metric as a means of comparing spectral functions, and to give an intuitive sense for what different values of this metric mean, we display in Figure 3 the spectral function of the molecule Limonene at high broadening and low field, computed for several different maximum cluster sizes, in each case comparing against the exact result (which for this molecule corresponds to a maximum cluster size of sixteen). In these plots, rather than plotting against the frequency directly, we plot against the quantity
$\Delta=\frac{\omega-\omega_{\text{ref}}}{\omega_{\text{ref}}}=\frac{\nu-\nu_{\text{ref}}}{\nu_{\text{ref}}},$ (13)
where $\omega_{\text{ref}}$ is, as usual, the Larmor frequency of TMS. In terms of this shifted and scaled variable, the spectral function has peaks located at the values of the chemical shifts $\delta$, rather than the Larmor frequencies themselves, which makes the identification of chemical groups simpler, and is the convention in the NMR literature (this is sometimes referred to as “plotting against the chemical shift”).
@@
With this understanding of the error metric in mind, we present results for the convergence of three selected molecules as a function of maximum cluster size, all of which contain sixteen nuclear spins. We display the high broadening case in Figure 1 and the low broadening case in Figure 2. As we would expect, the convergence generally improves for higher field and higher broadening, with the broadening seemingly having the more pronounced effect on accuracy. We highlight the fact that there are no sudden jumps in accuracy, either up or down, as a function of maximum cluster size, and that the error of the approximation is (almost always) decreasing. This fact will be important for attempting to benchmark the performance of the solver when applied to molecules for which we do not have an exact solution.
Figure 4: The molecule Triphenylphosphine oxide, with one phosphorous atom and fifteen hydrogen atoms.
However, it is true that the convergence for one of the molecules, Triphenylphosphine oxide (TPPO), is a bit more stubborn than the other two. This is especially egregious in the case of low broadening. An explanation for this behaviour can be found by examining the structure of the TPPO molecule, which can be seen in Figure 4. The active nuclei in this molecule are the 1H nuclei contained within the three benzene rings, along with a 31P nucleus connecting them, all of which leads to a highly symmetric spin Hamiltonian. However, as the three outermost 1H nuclei are not directly coupled to the 31P nucleus, the clusters chosen for these nuclei do not, by default, contain the 31P nucleus, since it is always ranked as the least important nucleus, and thus these clusters lack crucial information about the global structure of this highly symmetric molecule. It is therefore not surprising that the spectra which are computed using the basic clustering scheme do not converge well. We will have more to say about this case (as well as a related one) in a later section. There we will propose a refined clustering method that does lead to good convergence behaviour for such unusual cases, where a lack of direct couplings leads to clusters that conflict with the symmetry of the molecule.
In Appendix D we provide convergence data for additional selected molecules with nuclei counts of eight, nine, ten, and twelve. All of these molecules display strong convergence behaviour.
@@
V Larger Molecules
Here we discuss the performance of the solver when applied to larger molecules. Since we do not possess an exact solution for these molecules, the spectral functions for smaller cluster sizes are instead compared against the spectral function for the largest cluster size we can reasonably simulate (referred to as the “reference” size). As we examine the spectral functions corresponding to smaller cluster sizes, if we eventually reach a point beyond which we no longer see any significant differences between the smaller size and the reference size, then we will consider the solutions to have converged. While this does not represent a rigorous proof that we have converged to the correct solution, it is nevertheless plausible, given that we have not found any examples in the exactly solvable case in which the accuracy of the solution suddenly jumps from one approximate calculation to the next after having reached a plateau.
@@ V Larger Molecules
In Figures 7 and 8 we display the results of our analysis for a selection of ten molecules of various size, all of them with a reference cluster size of seventeen. Once again, we see that the convergence is better for higher field and higher broadening, and the agreement between a cluster size of sixteen and a cluster size of seventeen is exceptionally good in all cases. According to our error metric, the molecule with the worst convergence would appear to be Friedelin at low broadening, with error values of -2.77, -2.70, and -2.75 for high, low, and very low field, respectively. We also notice that the error values for the molecule tert-Butyl(tert-butylphosphanyl)phosphane (also named 1,2-di-tert-butyl-diphosphane, referred to here as “Diphosphane”) seem to progress in a somewhat erratic fashion from one cluster size to the next, for several parameter regimes. Even though these error values for Diphosphane are in principle very small, this erratic behaviour, along with the fact that Diphosphane is a highly symmetric molecule (similar to TPPO), is cause for suspicion. We briefly investigate the spectral functions of both of these molecules.
V.1 Friedelin
@@
In Figure 10 we display the spectral function of Friedelin for the case of low field and low broadening, which, according to our error metric, should represent the worst convergence of any spectral function. While there are some small details which are not perfectly captured, the majority of the larger features appear to be resolved, to the extent that it is not clear whether a more precise simulation would add to the practical interpretation of the NMR spectrum. For the case of Friedelin, this implies that our solver is capable of reaching the level of accuracy necessary for industrial applications, especially since this level of broadening is already somewhat experimentally unrealistic.
V.2 Diphosphane
Figure 11: The spectral function for Diphosphane, for the case of high field and low broadening using the simple clustering method. For the improved clustering see Fig. 16.
@@ V.2 Diphosphane
To understand the origin of this behaviour, in Figure 12 we show a schematic representation of the Diphosphane molecule. This molecule consists of two central phosphorous atoms, each of which is coupled to three Methyl groups that together contain nine hydrogen atoms. Additionally, each phosphorous atom has a hydrogen atom that is strongly coupled to it (leading to a total of twenty hydrogen atoms). The key feature of the NMR Hamiltonian for this molecule is that none of the hydrogen nuclei couple to each other directly. Rather, each hydrogen nucleus couples only to the nearest phosphorous nucleus. The Diphosphane molecule therefore represents another case in which a lack of direct couplings may lead to clustering which conflicts with the symmetry of the molecule, thereby leading to inaccurate results.
@@
Figure 13: The spectral function for Diphosphane, for the case of high field and low broadening, at larger cluster sizes, which we can obtain due to the symmetry of the molecule using the simple clustering method. For the improved clustering see Fig. 16.
@@
To investigate this possibility, we can in fact compare these results against the exact solution for Diphospane, due to its previously mentioned high level of symmetry. The two phosphorous nuclei have the same chemical shift as each other, which is also true for the two central hydrogen nuclei. These two central hydrogen nuclei couple with the same strength to their respective phosphorous nuclei. However, much more significant is the fact that all of the eighteen hydrogen nuclei contained in the Methyl groups have identical chemical shifts, and also couple identically to their respective phosphorous nuclei. Thus, all of the “magnetically equivalent” spins can be combined into a single (higher representation) spin object, by exploiting the local SU(2) symmetry. Specifically, this reduction leads to an effective system of two spin-9/2 particles coupled to four spin-1/2 particles, which can be solved in approximately 10 minutes of computational time on a laptop with 32 Gigabytes of memory and an Intel(R) Core(TM) Ultra 7 165U processor.
@@
In Figure 13 we again display the spectral function of Diphosphane at the two relevant portions of the spectrum, this time for larger cluster sizes. We continue to see the oscillating behaviour at higher chemical shifts, right up to the exact solution. In addition, we see a dramatic qualitative change in the spectrum at lower chemical shift. As we increase the cluster size from 21 to 22 spins (the full size of the system), a significant fraction of the weight in the two large peaks gets transferred to many other intermediate peaks which appear for the first time. This extreme behaviour verifies our suspicion that our basic clustering method is inadequate for this molecule.
These results we have seen for Diphosphane and TPPO clearly indicate that for molecules which exhibit a high level of symmetry and a small number of direct spin-spin interactions, a refinement to our basic clustering method must be introduced.
VI Improving the Clustering Scheme
Here we discuss a relatively simple extension to our basic clustering scheme which enables us to obtain the correct convergence behaviour for both Diphosphane and TPPO. We emphasize that our intent here is not to develop the best possible clustering method for these molecules. Rather, we simply wish to demonstrate that very simple improvements to the basic clustering scheme can already lead to a noticeable improvement for these molecules which possess very few direct couplings in the presence of high symmetry. An ideal clustering scheme might, for example, utilize higher order terms in the J couplings, or possibly invoke ideas from graph theory, and it should certainly include a careful consideration of all of the symmetries present in the structure of a given molecule. Such a study would, however, be outside the scope of this work.
With this caveat in mind, our extended scheme will construct the cluster for a given nucleus as follows. If the chosen cluster size does not exceed the number of other nuclei which couple directly to the given nucleus, then the clustering proceeds as usual. Otherwise, since we cannot fill the cluster with directly coupled nuclei, we turn to the nucleus which is most strongly coupled to the original given nucleus (according to the metric in eq. (10)), and designate it as the secondary nucleus. We now add further nuclei to the cluster, except that the coupling is now between the secondary nucleus and these additional nuclei. If the cluster is still not filled after considering all of the nuclei which are directly coupled to the secondary nucleus, then we find the nucleus which is most strongly coupled to either the first or second nuclei, and designate it as the tertiary nucleus, repeating the procedure with this third nucleus. This process repeats until the cluster has been filled.
Figure 14 displays the convergence behaviour of the three exactly solvable molecules we have previously considered, however now with the spectral functions obtained using this extended clustering. We choose to examine the most extreme case, at very low field and low broadening, in order to highlight the success of this extension. As we can see, the spectral function for TPPO now converges well, while the convergence behaviour of the other two molecules is not significantly altered.
Figure 14: The convergence for selected molecules with 16 nuclear spins for the case of low broadening and very low field, using the extended clustering.
A similar behaviour is reflected in the convergence properties of the larger molecules. Diphosphane is in fact the only one of these molecules which has its convergence behaviour impacted in any meaningful way by the extended clustering, and so we limit our discussion here to this molecule. Figure 15 displays the convergence for this molecule, which we note does indeed still demonstrate an unusual oscillating pattern. However, an examination of the spectral functions obtained with this new clustering, which we display in Figure 16, demonstrates that the proper qualitative features of the spectrum are now visible in the approximate spectral functions, with reasonably accurate results being seen when using clusters as small as eight spins. Moving beyond eight spins, the only changes we see in the spectrum concern the precise number of peaks in the individual multiplets, along with their exact position. However, these changes occur in a predictable fashion, as their positions alternate between even and odd sized clusters, with each additional spin leading to an additional peak splitting. As we have seen that this behaviour indeed continues throughout all larger cluster sizes, we come to the conclusion that a knowledge of the total number of spins in the molecule, together with the approximate results obtained using cluster sizes eight through eleven, would be sufficient to identify the correct spectrum of this molecule.
Figure 15: The convergence for Diphosphane for the case of high field and low broadening, using the extended clustering scheme.
Figure 16: The spectral function of Diphosphane, for the case of high field and low broadening, using the extended clustering scheme. The spectral functions for cluster sizes of eight through eleven already demonstrate good qualitative agreement with the exact solution at 22 spins.
Taken together, these results demonstrate that even for the most pernicious cases we have studied, only a minimal extension to our clustering scheme is necessary to obtain good convergence.
VII Conclusion
@@
In this work we have investigated the performance of our NMR solver when applied to a collection of selected molecules. Across all parameter regimes of practical relevance, we have found that our basic clustering scheme performs very well for almost every molecule, with the performance increasing with higher magnetic fields and higher broadening. The only exceptions take the form of highly symmetric molecules which exhibit very few direct spin couplings. However, a simple extension to our clustering scheme is sufficient to tackle even this case. This demonstrates the utility of our clustering-based solver as a powerful tool for understanding molecular NMR spectra, even for molecules that we can not solve exactly.
With this conclusion in mind, we turn to the question of potential quantum advantage in the context of NMR. At the relatively high magnetic fields which are utilized in conventional NMR experiments, the performance of our solver casts doubt on the possibility of demonstrating quantum advantage in the calculation of traditional proton NMR spectra. While we cannot rule out the possibility that there may exist sufficiently large, highly symmetric molecules of industrial relevance which may pose a challenge for our solver, one must also consider whether the number of such molecules would be sufficiently high as to justify the wide scale application of quantum computing in this context.
@@
At lower magnetic fields, however, the situation may be different. The decreasing performance of our solver with lower broadening and lower magnetic fields leads us to the domain of zero field NMR spectroscopy [2]. In such experiments, hyperpolarized samples are used to measure the NMR spectra at magnetic field strengths of only a few nanotesla, such that the Zeemann term can essentially be neglected. Furthermore, the broadening in these experiments can reach as low as 0.01 Hz, leading us to believe that classical solvers (including our own) should struggle in this regime. This therefore suggests a potential path towards quantum advantage [52].
The possibility of demonstrating quantum advantage in the context of zero field NMR is a question which we are currently investigating and hope to report on in more depth in the future.
@@ Appendix A Derivation of the Spectral Function
$C\left(\omega\right)\propto\eta\sum_{n,m}\frac{\langle E_{n}|\hat{M}^{-}|E_{m}\rangle\langle E_{m}|\hat{M}^{+}|E_{n}\rangle}{\eta^{2}+[\omega-(E_{n}-E_{m})]^{2}}$ (14)
is connected to measurements made in an NMR experiment, specifically the so-called pulse-acquire experiment. To this end, we will first discuss what properties are measured in such an experiment, and then derive from them the spectral function.
@@ A.1 The pulse acquire experiment
In an NMR experiment, the magnetic field resulting from the magnetization of the sample is measured. The average magnetic moment of a single molecule in the sample along a specified direction is given by the expectation value of the appropriate spin operator,
$m^{\alpha}=\hbar\sum_{i}\gamma_{i}\langle\hat{I}^{\alpha}\rangle=\hbar\langle\hat{M}^{\alpha}\rangle$ (15)
@@
$\hat{M}^{\alpha}\equiv\sum_{i}\gamma_{i}\hat{I}_{i}^{\alpha},$ (16)
with $\alpha\in\{x,y,z\}$. To connect this object to the magnetic field measured in an NMR experiment, we note three important facts. First, the induced magnetic field due to a single molecule is proportional to its magnetic moment. Second, a measurement of the total magnetic field is essentially an ensemble average over macroscopically many quantum measurements, one for each molecule in the sample, and is thus proportional to the expectation value above. Third, since we are not worried about the overall normalization of the spectral function, any overall proportionality factors, such as the permeability or the total number of molecules, are irrelevant. For these reasons, we can simply study the expectation value of the magnetic moment of a single molecule, and our task is therefore to understand how this quantity evolves in time in an NMR experiment.
@@
After a sample is placed in an NMR spectrometer, it is allowed to come to thermal equilibrium, a state described by the density matrix
$\rho_{\beta}=\frac{1}{Z}\mathrm{e}^{-\beta\hbar\hat{\cal{H}}}$. (17)
@@
with $Z$ the partition sum. For a spectrometer approaching the limits of current technology, with a natural frequency $\nu$ of 1 GHz, the characteristic energy scale of the molecular NMR Hamiltonian is
$h\nu\approx 6.626\times 10^{-25}~\text{J},$ (18)
while the thermal energy scale at a temperature of only one Kelvin is
$k_{B}T\approx 1.381\times 10^{-23}~\text{J}.$ (19)
@@
Since the thermal energy scale is several orders of magnitude larger than the characteristic energy scale of the molecule in any conceivable scenario, it is an incredibly good approximation to take
@@
$\begin{split}Z\rho_{\beta}&=\mathds{1}-\beta\hbar\hat{\cal{H}}\\ &=\mathds{1}-\beta\hbar\left[-\sum_{l}\gamma_{l}B^{z}\hat{I}^{z}_{l}+2\pi\sum_{k<l}J_{kl}\mathbf{\hat{I}}_{k}\cdot\mathbf{\hat{I}}_{l}\right],\end{split}$ (20)
@@
where we have again ignored the correction due to the chemical shifts. Furthermore, even at magnetic fields as low as the Earth’s magnetic field, the coupling terms are several orders of magnitude smaller than the Zeeman terms, and thus these operators can be ignored in the thermal density matrix above. We thus take
$Z\rho_{\beta}=\mathds{1}+\beta\hbar B^{z}\sum_{l}\gamma_{l}\hat{I}^{z}_{l}=\mathds{1}+\beta\hbar B^{z}\hat{M}^{z}$ (21)
Once the system has reached its equilibrium state, an electromagnetic pulse is applied, which is implemented by sending radiofrequency current through a coil wrapped around the sample. Through proper control of the phase, frequency, amplitude, and duration of this pulse, it is possible to manipulate the nuclear spins in the sample, and depending on the frequency band where this pulse is effective, different isotopes can be excited. In the standard proton NMR pulse-acquire experiment which we discuss here, a single pulse close to the Larmor frequency of TMS is used to excite the hydrogen nuclei, and we assume that it is possible to engineer this pulse such that all hydrogen nuclei are affected equally. If we view the system from a frame of reference that rotates with the Larmor frequency of TMS, we can describe the effect of the pulse as inducing a $\pi/2$ rotation of the net magnetization of these nuclei into the XY plane (a rotation which we take to be around the X axis). The hydrogen portion of the density matrix in this frame is therefore transformed as
$\rho_{\text{Hy}}\to\exp\left(-i\frac{\pi}{2}\hat{I}^{x}\right)\rho_{\text{Hy}}\exp\left(+i\frac{\pi}{2}\hat{I}^{x}\right).$ (22)
If only hydrogen nuclei are present in the molecule, then applying this transformation to the thermal density matrix (which takes the same form in the rotating frame) yields
$Z\rho_{\beta}\to Z\rho_{0}=\mathds{1}-\beta\hbar B^{z}\hat{M}^{y}$ (23)
as the initial density matrix after the pulse. Contributions from other nuclei in the molecule, which are not flipped by the initial pulse, will eventually drop out of the formula for the net magnetization (for symmetry reasons similar to the one that allows us to ignore the identity operator later on). As a result, we can simply omit any contributions from other nuclei from this point on.
After the pulse has been applied, the system is allowed to evolve freely in time, and the resulting magnetization is measured. It is convenient to continue working in the rotating frame of reference, with the density matrix of the rotating frame and the density matrix of the laboratory frame being related by
$\rho_{L}=\mathrm{e}^{-i\hat{\cal{H}}_{0}t}~\rho_{R}~\mathrm{e}^{+i\hat{\cal{H}}_{0}t},$ (24)
with
$\hat{\cal{H}}_{0}=-\omega_{\text{ref}}\sum_{l}\hat{I}^{z}_{l}\equiv-\omega_{\text{ref}}\hat{I}^{z}$ (25)
corresponding to uniform precession at the Larmor frequency $\omega_{\text{ref}}$ of the hydrogen nuclei in the reference compound (TMS for proton NMR). The rotating frame density matrix evolves according to the shifted Hamiltonian
$\hat{\cal{H}}_{R}=\hat{\cal{H}}-\hat{\cal{H}}_{0}=-\sum_{l}\left(\omega_{l}-\omega_{\text{ref}}\right)\hat{I}^{z}_{l}+2\pi\sum_{k<l}J_{kl}\mathbf{\hat{I}}_{k}\cdot\mathbf{\hat{I}}_{l}.$ (26)
We note that for a hydrogen nucleus in particular, the definition of the chemical shifts leads to
$\omega_{l}-\omega_{\text{ref}}=\gamma_{\text{ref}}(1+\delta_{l})B^{z}-\gamma_{\text{ref}}B^{z}=\gamma_{\text{ref}}\delta_{l}B^{z},$ (27)
where $\gamma_{\text{ref}}$ is the gyromagnetic ratio with respect to which the chemical shifts have been defined. The density matrix of the molecule in the rotating frame therefore evolves in time according to
$\rho\left(t\right)=\mathrm{e}^{-i\hat{\cal{H}}_{R}t}\rho_{0}\mathrm{e}^{+i\hat{\cal{H}}_{R}t}\propto\mathds{1}-\beta\hbar B^{z}\mathrm{e}^{-i\hat{\cal{H}}_{R}t}\hat{M}^{y}\mathrm{e}^{+i\hat{\cal{H}}_{R}t}.$ (28)
@@
$m^{\alpha}=\hbar\langle\hat{M}^{\alpha}\rangle=\hbar\text{Tr}\left[\hat{M}^{\alpha}\rho\left(t\right)\right].$ (29)
We can, if we like, omit nuclei other than hydrogen from $\hat{M}^{x/y}$ in the above expression, since these do not yield a significant signal in a proton NMR spectrum (this can be shown via a perturbative analysis). Since $\hat{M}^{\alpha}$ is traceless, only the second term in the density matrix contributes to the trace, and we have
$m^{x}=\hbar\langle\hat{M}^{x}\rangle\propto\text{Tr}\left[\hat{M}^{x}\mathrm{e}^{-i\hat{\cal{H}}_{R}t}\hat{M}^{y}\mathrm{e}^{+i\hat{\cal{H}}_{R}t}\right],$ (30)
@@
$m^{y}=\hbar\langle\hat{M}^{x}\rangle\propto\text{Tr}\left[\hat{M}^{y}\mathrm{e}^{-i\hat{\cal{H}}_{R}t}\hat{M}^{y}\mathrm{e}^{+i\hat{\cal{H}}_{R}t}\right],$ (31)
@@
$m^{\alpha}\to\mathrm{e}^{-\eta t}m^{\alpha}\,.$ (32)